On June 30, 2021, Georgia-Atlantic, Inc. leased a warehouse equipment from IC Leasing Corporation. The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $681,229 over a four-year lease term, payable each June 30 and December 31, with the first payment at June 30, 2021. Georgia-Atlantic's incremental borrowing rate is 8%, the same rate IC uses to calculate lease payment amounts. Amortization is recorded on a straight-line basis at the end of each fiscal year. The fair value of the equipment is $4.8 million. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Determine the present value of the lease payments at June 30, 2021 that Georgia-Atlantic uses to record the right-of-use asset and lease liability.
2. What pretax amounts related to the lease would Georgia-Atlantic report in its balance sheet at December 31, 2021?
3. What pretax amounts related to the lease would Georgia-Atlantic report in its income statement for the year ended December 31, 2021?
(For all requirements, enter your answers in whole dollars and not in millions. Round your final answers to the nearest whole dollar.)
Answer #1

Solution 1:

Semiannual lease payment = $681,229

Total semiannual payments = 4*2 = 8

Incremental borrowing rate = 8%, 4% semiannual

Present value of minimum lease payments used to record right to use assets = Semi Annual lease payments * Cumulative PV Factor of annuity due for 8 periods at 4%

= $681,229 * 7.00205 = $4,770,000

Solution 2:

Semiannual payment on 30.06.2021 = $681,229

Pretax amount of liability on 30.06.2021 = ($4,770,000 - $681,229) = $4,088,771

Interest expense for 31.12.2021 = $4,088,771 * 4% = $163,551

Semiannual lease payment on 31.12.2021 = $681,229

Pre tax amount for liability December 31, 2021 = $4,088,771 + $163 ,551- $681,229 = $3,571,093

Depreciation on right to use assets for 2021 = $4,770,000 / 4 * 6/12 = $596,250

Pre tax amount of right to use asset to be reported for 2021 = $4,770,000 - $596,250 = $4,173,750

Solution 3:

Pre tax amount of interest expense Georgia Atlantic Inc. reports in its income statement = $4,088,771 * 4% = $163,551

Pre tax amount of amortization expenses Georgia Atlantic Inc. reports in its income statement = $4,770,000 / 4 * 6/12 = $596,250
Answer #2

page 1 Given data that On june 30 2021 Geosigia Atlantic Inc leased a wagle house equipment from lc leasing compogatio The fapage 2 + BT в обеаѕе payments 681229 a dease Tego B2*2 3. Bogelowing fate 8/100 B32 4 Fais value of equipment 4800000 5 CumulPage 3 B 681,229 4 - 8% 4,800,000 4.1 Lease payment 12 Gease Team 13 Bosilowing Rate y Fairy Value of Equipment 5 Cumulative

answered by: pro